OpenStudy (anonymous):

What value of k will make the equation below true?7/4k-9=3 a.3 3/7 B. 6 6/7 c. 12 D 21

OpenStudy (anonymous):

do you mean\[7/(4k-9)=3\] ?

OpenStudy (anonymous):

4k-9=3 solve for k

OpenStudy (anonymous):

\[7/(4k-9)=3/1\]

OpenStudy (anonymous):

now cross multiply

OpenStudy (anonymous):

Not sure I know the answer is21 but I don't know how to get there. Sorry that is wrong the answer is 6 /67 When I solve for K I get 3 and that is wrong. I am horrible at this Then how is the answer 6 6/7 six and 6 sevenths. I have the answer key and I have to figure out how to get the answer as I was wrong.

OpenStudy (anonymous):

Oh 33/12 = 11/4 I have to get it in 7ths

OpenStudy (anonymous):

This is 7th grade pre algebra

OpenStudy (anonymous):